### v2

# SciReplicate-Bench: Benchmarking LLMs in Agent-driven Algorithmic Reproduction from Research Papers

Abstract
[…]
1 Introduction
The evolution of Large Language Models (LLMs) has ushered in a transformative era in scientific discovery, positioning them as powerful tools for streamlining research (Gridach et al., 2025; Buehler, 2024; Lu et al., 2024), from idea generation to verification and publication writing. For instance, Si et al. (2025); Gu & Krenn (2025) demonstrated how LLMs can be prompted to generate novel research ideas, while Yuan et al. (2022); Du et al. (2024) explored their use in producing literature reviews for idea evaluation. Additionally, LLMs are increasingly integrated into tools like Semantic Scholar [^1] https://www.semanticscholar.org, Research Rabbit [^2] https://www.researchrabbit.ai/, and Undermind Research Assistant [^3] https://www.undermind.ai/, enhancing literature discovery, citation analysis, and knowledge synthesis. These advancements, both in research methodologies and practical applications, suggest that LLMs have the potential to assist across multiple stages of scientific discovery.
Among the aforementioned advancements in research acceleration, the ability of LLMs to correctly generate code for validating real-world scientific ideas is particularly noteworthy. Computational validation is crucial across many fields, yet researchers often face barriers due to limited coding expertise or inaccessible implementations. By converting scientific algorithm descriptions into executable code, LLMs could enhance reproducibility and accelerate scientific discovery. However, despite progress in LLM-based code generation, a significant gap remains in generating code directly from scholarly papers. First, algorithm comprehension from scientific papers is challenging. Research papers are characterized by their brevity, methodological rigor, and extensive citations, with critical details about algorithms often dispersed across multiple sections of the paper. Understanding these algorithms requires synthesizing information from internal references and external scholarly works. Second, code repositories typically consist of multiple interdependent files and directories. To implement an algorithm, LLMs must comprehensively examine file dependencies, identify reusable components, and correctly handle both internal dependencies and external APIs.
Despite the importance of automated scientific idea verification, there exists no dataset specifically designed to evaluate the ability of LLMs to reproduce real-world algorithms proposed in peer-reviewed publications. As shown in Table 1, there are several machine learning software engineering benchmarks primarily focused on evaluating algorithmic design or straightforward implementations, which are significantly less complex than the methods typically described in academic research papers. For example, MLE-BENCH (Liu et al., 2023) and MLAgentBench (Huang et al., 2023) utilize Kaggle competitions, where LLMs must develop and implement solutions based on provided task specifications. ML-BENCH (Chan et al., 2024) uses Machine Learning (ML) GitHub repositories to assess LLMs’ text-to-code capabilities and test autonomous agents in task execution.
Benchmark Paper Understanding Repo-Search Test Case Source Task Types
MLE-BENCH Kaggle Algorithm design and code gen.
MLAgentBench Kaggle Algorithm design and code gen.
ML-BENCH Github Code gen.
SciReplicate-Bench Publications Replicate code for algorithms
in real-world NLP publications.
Table 1: Comparisons of different machine learning software engineering benchmarks.
To address this gap, we developed SciReplicate-Bench, the first benchmark specifically designed to evaluate LLMs’ capabilities in code generation for reproducing research findings from academic papers. It consists of 100 code reproduction tasks derived from 36 papers published in leading conferences in 2024. This recent publication window was deliberately chosen to minimize the risk of data leakage. An overview of the task is illustrated in Figure 1, with a concrete example provided in Figure A2 in Appendix E. The task consists of two main steps: 1. Algorithm understanding. LLMs must extract essential information from the paper, such as workflow details, algorithm descriptions, and hyperparameter values. 2. Code implementation. LLMs then implement a function or method within a provided repository, using both the extracted information and the LaTeX representation of the algorithm from the paper. We introduce Sci-Reproducer, a dual-agent system that combines a Paper Agent and a Code Agent to handle these two steps collaboratively and implement code for the target algorithm.
To rigorously assess LLM performance on this benchmark, we evaluate two dimensions corresponding to the aforementioned two steps: algorithm comprehension correctness and code correctness. To evaluate algorithm comprehension, we introduce a reasoning graph to represent the reasoning logic behind algorithm reproduction. Each node in the graph represents a code comment, which reflects a single reasoning step and is aligned with a specific segment of code. Edges between nodes are defined based on data flow relationships across different code segments. We compute the similarity between the generated reasoning graph and a reference graph to derive the reasoning graph accuracy. To evaluate code correctness, we employ established metrics including execution accuracy (Rajkumar et al., 2022; Xiang et al., 2023), CodeBLEU (Ren et al., 2020), and recall of intra/cross-file dependencies and APIs.
Our work makes the following contributions:
  • Benchmarks: SciReplicate-Bench, a benchmark of 100 algorithm reproduction tasks from recent NLP publications.
  • Metric: We propose a novel reasoning graph accuracy metric for evaluating algorithmic comprehension.
  • Approach: Sci-Reproducer, a dual-agent framework combining paper understanding and code implementation.
  • Insights: Comprehensive evaluation across state-of-the-art LLMs reveals four key findings: (i) the task remains highly challenging, with execution accuracy below 40% for all models; (ii) reasoning models exhibit “overthinking” behavior (Cuadron et al., 2025; Sui et al., 2025), over-relying on internal reasoning rather than utilizing available tools for information extraction; (iii) while LLMs demonstrate strong algorithmic comprehension, they struggle with practical implementation; and (iv) algorithm reproduction failures often stem from incomplete paper descriptions, which our Sci-Reproducer effectively addresses.
2 Related Work
Our work lies at the intersection of AI for automating scientific discovery and LLM-based code generation.
2.1 AI for Automating Scientific Discovery
The application of LLMs to accelerate scientific research has emerged as a rapidly growing field with diverse approaches. Several studies have demonstrated the potential for comprehensive research automation through end-to-end AI systems. Schmidgall et al. (2025); Lu et al. (2024) developed frameworks that integrate idea generation, experimental validation, and manuscript composition, with some AI-authored papers successfully passing workshop review processes (Yamada et al., 2025). Complementary research has focused on the creative aspects of scientific inquiry, with Wang et al. (2023); Ghafarollahi & Buehler (2024); O’Neill et al. (2025) investigating LLMs’ capacity for generating novel research hypotheses. Notably, recent evaluations by Gu et al. (2024); Kumar et al. (2024); Liu et al. (2025); Si et al. (2024) suggest that AI-generated research concepts may occasionally exceed human-generated ideas in terms of novelty and originality.
Within computational disciplines where implementation validation is essential, LLMs have shown promise in algorithm design and code development tasks. Our proposed SciReplicate-Bench addresses a previously underexplored area: the automated reproduction of algorithms directly from academic publications. This represents a unique challenge at the intersection of scientific literature comprehension and executable code synthesis. While recent parallel efforts such as PaperBench (Starace et al., 2025) and Paper2CodeBench (Seo et al., 2025) have tackled related problems by exploring full codebase reconstruction, these evaluation approaches rely substantially on manual assessment criteria and LLM-based correctness judgments, introducing potential inconsistencies and reliability concerns. Our approach prioritizes objective evaluation through execution accuracy, providing more rigorous validation than non-executable assessment methodologies (Wang et al., 2022; Chen et al., 2021).
The broader ecosystem of computational reproducibility research includes specialized frameworks such as MLGym (Nathani et al., 2025) for baseline improvement, and evaluation benchmarks developed by Siegel et al. (2024); Ren et al. (2023) that assess LLMs’ ability to reproduce published experimental results using existing codebases.
2.2 LLMs for Code Generation
Code generation has emerged as a prominent application of LLMs, with benchmarks ranging from basic programming tasks (Chen et al., 2021; Jain et al., 2024; Austin et al., 2021; Hendrycks et al., 2021; Liu et al., 2022) to realistic software engineering challenges like SWE-bench (Jimenez et al., 2023), which uses actual repository pull requests. However, these benchmarks primarily target general software engineering rather than scientific algorithm reproduction.
Recent efforts have developed machine learning-specific benchmarks (Liu et al., 2023; Huang et al., 2023; Chan et al., 2024), but these typically involve implementing algorithms proposed by the models themselves or solving relatively straightforward tasks. They lack the depth of algorithmic understanding and rigorous paper analysis required for reproducing algorithms from peer-reviewed publications.
Despite advances in tool-augmented code generation (Schick et al., 2023; Zhang et al., 2024a; a; 2023b), no existing system specifically addresses the unique challenge of translating academic papers into executable code. Our Sci-Reproducer framework demonstrates the ability to comprehend academic publications and convert abstract algorithm descriptions into functional implementations.
3 SciReplicate-Bench
Overview
SciReplicate-Bench is designed to evaluate LLMs’ ability to reproduce algorithms from academic papers, consisting of 100 tasks curated from 36 recent NLP publications with their corresponding open-source implementations. The task categories are detailed in Figure A1 in Appendix B. The benchmark focuses on repository-level code generation, where each task is centered around implementing a specific function or class method. As illustrated in Figure A3 in Appendix E, each task comprises nine components, which can be categorized into three groups corresponding to code generation, evaluation, and analysis, respectively.
For code generation, the following components are provided as inputs to LLMs:
  1. Function signature: the definition of the target function, including detailed descriptions of its input and output variables.
  2. Algorithm Description: The LaTeX code description of the target algorithm, typically located within a subsection or paragraph of the target paper.
  3. Literature context: the original paper along with its cited references, providing broader conceptual context.
  4. Repository context: all source files and code in the repository that inform or support the target implementation.
For evaluation, the following components are provided for code execution and metrics calculation:
  1. Reference implementation: ground-truth code serving as the reference for CodeBLEU evaluation.
  2. Reasoning graph annotations: structured representations of the algorithmic logic and implementation flow, enabling assessment of reasoning graph accuracy.
  3. Dependency annotations: comprehensive documentation of internal dependencies, cross-file relationships, and external API usage for computing recall metrics across all dependency categories.
  4. Test environment: isolated Python execution environment containing validation cases and automated verification scripts for assessing implementation correctness.
To enable further analysis of the underlying causes of LLM failures, the benchmark includes:
  1. Missing/Mismatch Information: the LaTeX description of the algorithm may omit certain implementation details, which could either appear elsewhere in the paper or be entirely absent. We also annotate mismatches between the paper description and the reference implementation.
Task Definition
Based on SciReplicate-Bench, an LLM is given the algorithm description, function signature, literature context, and repository context as input. The LLM is asked to output a function that implements the target algorithm.
Benchmark Construction
The benchmark construction process comprises four key steps: paper selection, Python environment setup, documentation, and verification suite preparation. To mitigate the risk of data leakage, we selected papers published in 2024 that provide publicly available code repositories. During the annotation process, each repository was refactored to isolate the core algorithm into a standalone function, and all sources of randomness were removed to ensure reproducibility and prevent leakage. On average, annotating each paper requires approximately 12 hours. Details of the annotation process are provided in Appendix A1.
3.1 Evaluation Metrics
3.1.1 Evaluating Algorithm Comprehension
We propose the reasoning graph accuracy metric to evaluate how well LLMs understand the logic and implementation of algorithms. During code generation, LLMs are prompted to insert specially formatted, non-overlapping, non-nested comments that mark reasoning steps derived from the algorithm’s LaTeX code (The prompt can be found in Figure A5). We then construct a reasoning graph $G=\{N,E\}$ (illustrated in Figure A3), modeled as a Directed Acyclic Graph (DAG). Each node $n_{i}=\langle w_{i},c_{i}\rangle,n_{i}\in N$ represents a reasoning step with a comment $w_{i}$ and corresponding code snippet $c_{i}$. An edge $e_{i}=\langle n_{i},n_{j}\rangle,e_{i}\in E$ is added if a variable used in $c_{j}$ is defined or last modified in $c_{i}$. To compute the reasoning graph accuracy, we compare the generated graph $G_{g}$ with the reference graph $G_{r}$ via node and edge matching:
  1. Node matching: comments from $G_{r}$ and $G_{g}$ are passed to GPT-4o, which maps each reference node to one or more nodes in the generated graph. A node in $G_{r}$ is considered matched if it has at least one corresponding node in $G_{g}$. The prompt template used for this process is available in Figure A4.
  2. Edge matching: for each reference edge $e_{r}=\langle n_{r}^{i},n_{r}^{j}\rangle$, if both endpoint nodes have corresponding matches in $G_{g}$, we apply Breadth-First Search(BFS) to verify whether a corresponding edge exists in $G_{g}$.
The reasoning graph accuracy $S_{r}$ is computed as:
$S_{r}=\sum_{n_{i}}^{n_{i}\in N_{m}}s^{n}_{i}+\sum_{e_{j}}^{e_{j}\in E_{m}}s^{e}_{j}.$ (1)
where $N_{m}$ and $E_{m}$ denote the sets of matched nodes and edges, respectively, and $s_{i}^{n}$ and $s_{j}^{e}$ represent their corresponding significance scores. Node significance is determined by the complexity of its corresponding code segment, measured by the number of variable definitions and usages, function calls, arithmetic operations, and lines of code, then normalized across the reference graph. Edge significance is calculated as the product of the significance scores of its connected nodes, followed by normalization.
3.1.2 Evaluating Code Generation
For assessing coding ability, we use the following evaluation metrics:
  • Execution accuracy (Xiang et al., 2023; Zhang et al., 2024b; Long et al., 2022): we integrate the generated code into the repository and execute it to obtain results. If all test cases match the reference results, we consider the code correct.
  • CodeBLEU (Ren et al., 2020): this metric evaluates how similar the generated code is to reference code by using the traditional BLEU metric (Papineni et al., 2002) while incorporating syntactic information through abstract syntax trees (AST) and semantic understanding via data-flow graphs (DFG).
  • Recall (Li et al., 2024): we calculate recall scores specifically for intra-file dependencies, cross-file dependencies, and external APIs.
4 Sci-Reproducer
Figure 1: Overview of the task and the proposed Sci-Reproducer framework. The task involves algorithm understanding and code implementation, handled by a Paper Agent and a Code Agent operating in separate contexts with specialized actions.
Action Name Input Observation
Paper Agent
SearchPaper Query The retrieved response from the target paper in relation to
the query.
SearchSection Section ID The entire content of a section based on the section label.
SearchLiterature Paper ID, query The answer to the query searched from the literature (identified
by Paper ID).
Code Agent
SearchCode Name The definition of a specific code element in repository.
SearchFile Name The content of a certain file in repository.
SearchWeb Query The information obtained from the website.
Compiler code The feedback from the compiler after executing the code.
Table 2: The pre-defined actions for the Paper Agent and the Code Agent.
To address this task, we introduce Sci-Reproducer  [^4] A video demonstration showcasing Sci-Reproducer’s capabilities is provided at https://youtu.be/qcSIMgyehjE, a dual-agent framework designed for scientific paper algorithm replication. As illustrated in Figure 1, Sci-Reproducer  comprises a Paper Agent and a Code Agent that collaboratively work to replicate algorithms described in a given paper. The predefined actions employed by the agents are summarized in Table 2, with implementation details provided in Appendix C.
4.1 Paper Agent
Based on the provided algorithm description, the Paper Agent systematically retrieves contextual information from the literature context to support algorithmic understanding. Due to the input length limitations of LLMs, it is infeasible to input entire paper along with their associated literature. Consequently, the Paper Agent must selectively extract pertinent information, following a strategy akin to Retrieval Augmented Generation (RAG) (Wang et al., 2024b; Sarthi et al., 2024). The Paper Agent incrementally builds an understanding of the target algorithm by executing predefined actions to query the literature context. To facilitate this process, we adopt ReAct (Yao et al., 2022) as the agent strategy, which enables seamless integration of action execution with intermediate reasoning steps.
After the Paper Agent concludes that all necessary information has been collected, it generates a comprehensive literature report comprising key findings that fill in the missing components of the target algorithm’s LaTeX source. An example of the literature report is shown in Figure A8. This report subsequently serves as a crucial input for the Code Agent. The prompt used to guide the Paper Agent is provided in Figure A6.
4.2 Code Agent
Informed by the algorithm description, literature report, and code context, the Code Agent searches the repository to locate essential dependencies required for implementation. It can also browse websites for additional information and use a compiler to test and iteratively debug the code, ensuring proper execution by identifying and fixing syntax errors. The prompt for the Code Agent is provided in Figure A7.
5 Experiments
Approach Exe Acc($\uparrow$) CodeBLEU($\uparrow$) RG Acc($\uparrow$) Recall
Intra-File($\uparrow$) Cross-File($\uparrow$) API($\uparrow$)
[Uncaptioned image] GPT-4o-mini
No Agent 0.030 0.238 0.694 0.012 0.000 0.217
Paper Agent 0.040 0.246 0.717 0.024 0.000 0.251
Code Agent 0.140 0.279 0.738 0.565 0.364 0.328
Sci-Reproducer 0.170 0.303 0.741 0.576 0.364 0.362
[Uncaptioned image] GPT-4o
No Agent 0.040 0.259 0.705 0.059 0.000 0.281
Paper Agent 0.020 0.263 0.724 0.023 0.000 0.298
Code Agent 0.260 0.325 0.748 0.682 0.576 0.421
Sci-Reproducer 0.270 0.329 0.751 0.688 0.636 0.417
[Uncaptioned image] Claude-Sonnet-3.7
No Agent 0.070 0.282 0.725 0.094 0.091 0.362
Paper Agent 0.050 0.291 0.727 0.082 0.091 0.379
Code Agent 0.320 0.394 0.764 0.765 0.545 0.545
Sci-Reproducer 0.390 0.401 0.773 0.776 0.636 0.626
[Uncaptioned image] Gemini-2.0-Flash
No Agent 0.070 0.275 0.688 0.071 0.000 0.294
Paper Agent 0.040 0.278 0.699 0.082 0.000 0.332
Code Agent 0.220 0.323 0.708 0.553 0.212 0.426
Sci-Reproducer 0.250 0.346 0.727 0.588 0.333 0.455
[Uncaptioned image] Deepseek-V3
No Agent 0.030 0.260 0.712 0.012 0.061 0.272
Paper Agent 0.050 0.275 0.732 0.012 0.030 0.306
Code Agent 0.210 0.312 0.736 0.482 0.182 0.383
Sci-Reproducer 0.220 0.334 0.738 0.565 0.333 0.443
[Uncaptioned image] o3-mini-low [Uncaptioned image]
No Agent 0.080 0.259 0.758 0.035 0.000 0.323
Paper Agent 0.050 0.262 0.729 0.035 0.000 0.315
Code Agent 0.150 0.278 0.803 0.306 0.000 0.348
Sci-Reproducer 0.180 0.280 0.771 0.376 0.121 0.328
[Uncaptioned image] o3-mini-medium [Uncaptioned image]
No Agent 0.040 0.263 0.729 0.035 0.000 0.336
Paper Agent 0.060 0.263 0.726 0.047 0.000 0.319
Code Agent 0.220 0.289 0.749 0.376 0.030 0.404
Sci-Reproducer 0.240 0.283 0.758 0.341 0.061 0.362
[Uncaptioned image] o3-mini-high [Uncaptioned image]
No Agent 0.070 0.269 0.718 0.047 0.000 0.345
Paper Agent 0.070 0.267 0.751 0.035 0.000 0.366
Code Agent 0.160 0.277 0.773 0.165 0.152 0.374
Sci-Reproducer 0.160 0.283 0.763 0.294 0.091 0.357
Table 3: Performance evaluation on the SciReplicate-Bench benchmark. Models with [Uncaptioned image] notation indicate reasoning LLMs. “Exe Acc” represents execution accuracy while “RG Acc” indicates reasoning graph accuracy.
We evaluate Sci-Reproducer on the SciReplicate-Bench  benchmark using 7 advanced LLMs, including five non-reasoning LLMs: GPT-4o-mini (4o mini, 2024), GPT-4o (GPT-4o, 2024), Claude-Sonnet-3.7 (Claude-Sonnet-3.7, 2025), Gemini-2.0-Flash (Gemini-2.0-Flash, 2024), and Deepseek-V3 (DeepSeek-AI et al., 2024), and different versions of the reasoning models O3-mini (o3 mini, 2024), i.e., three different levels of reasoning intensity. For the reasoning graph accuracy metric, node matching is performed using GPT-4o, which may introduce some randomness. To reduce this variability, we set the temperature to 0 and top-p to 1, ensuring more deterministic generation. The calculation is repeated three times, and we report the average score as the final result.
5.1 Results on SciReplicate-Bench
Table 3 displays Sci-Reproducer’s evaluation results and contributions of Code/Paper Agent. The “No Agent”directly prompts the LLM to generate code based solely on the algorithm description and function signature. “No Paper Agent” allows the LLM to use Code Agent actions, but restricts access to Paper Agent actions. “No Code Agent” grants access to Paper Agent actions but blocks Code Agent capabilities. The results offer key insights, discussed in the following.
LLMs struggles on SciReplicate-Bench Most LLMs perform poorly, achieving less than 0.1 execution accuracy without using the agent to examine literature and repository contexts. With enhancement of Sci-Reproducer, these LLMs show notable improvements, with an average increase of 0.181 in execution ACC and 0.057 in CodeBLEU, although even the best-performing model, Claude-Sonnet-3.7, only achieved 0.390 execution accuracy. This highlights the exceptional challenge presented by our SciReplicate-Bench.
LLMs can comprehend algorithm logic LLMs demonstrate strong algorithmic comprehension capabilities, as evidenced by reasoning graph accuracy scores averaging 0.716 even without agent assistance. The addition of individual agents provides modest but consistent improvements: the Paper Agent increases understanding by 0.009 on average, while the Code Agent contributes a larger gain of 0.036. Combined agent deployment yields a cumulative improvement of 0.037. These enhancements stem from complementary mechanisms: the Paper Agent strengthens theoretical comprehension by gathering relevant contextual information from academic literature, while the Code Agent facilitates practical understanding through extraction of pertinent code patterns and dependency structures from repositories.
LLMs face challenges with actual implementation Although LLMs are capable of understanding algorithms, their performance in code generation remains suboptimal. Despite using Sci-Reproducer, the average execution accuracy remains low at 0.235, with a CodeBLEU score of 0.320.
Accurate dependency and API identification is crucial for code implementation Effectively recognizing and leveraging dependencies from the source repository and external APIs is essential for accurate code implementation. The integration of Code Agent led to substantial gains in recall with average increases of 0.441, 0.239, and 0.100, respectively, compared to cases without the agent. With Sci-Reproducer, Claude-Sonnet-3.7 attains the highest execution accuracy of 0.390, with the highest recall for intra/cross file dependency and API usage, at 0.776, 0.636, and 0.626 respectively.
Overthinking leads to limited improvement in reasoning LLMs Reasoning LLMs exhibit more modest performance gains when using Sci-Reproducer. While they achieve an average execution accuracy improvement of 0.13, non-reasoning models demonstrate substantially larger gains of 0.212. This pattern extends to recall metrics, where reasoning LLMs show improvements of 0.243, 0.061, and 0.041 respectively, compared to non-reasoning LLMs’ more pronounced gains of 0.560, 0.345, and 0.135. We attribute this performance gap to the “overthinking” phenomenon (Cuadron et al., 2025; Sui et al., 2025), where excessive internal reasoning impedes effective action execution, a limitation that we examine in detail in the following subsection.
5.2 Tool Usage Analysis
Figure 2: A grouped bar chart illustrating the frequency of tool usage by different models. The x-axis represents various actions, while the y-axis indicates the total number of times each tool was used on this dataset.
Figure 2 presents the number of times each LLM invokes actions on the full dataset with Sci-Reproducer. We observe the following:
  • For code-related actions, reasoning LLMs demonstrate limited tool usage, employing “SearchFile”, “SearchCodeItem”, and “SearchWeb” only 25.0, 3.3, and 0.0 times on average, respectively. Non-reasoning LLMs use these same actions far more extensively, with averages of 210.4, 68.2, and 16.8 times respectively. This disparity reveals a fundamental behavioral difference: reasoning models favor internal deliberation over external information gathering. Conversely, reasoning LLMs invoke “Compiler” more frequently, indicating they require more debugging iterations due to inadequate contextual information gathering. This over-reliance on internal reasoning undermines performance: advanced models like o3-mini-high and o3-mini-low achieve execution accuracy comparable to GPT-4o-mini, negating their theoretical computational advantages.
  • Paper-related actions exhibit a similar pattern. Reasoning LLMs use “SearchPaper”, “SearchSection”, and “SearchLiterature” an average of 56.3, 70.0, and 20.0 times respectively, while non-reasoning LLMs demonstrate substantially higher usage at 244.8, 188.4, and 58.0 times respectively. Additionally, we observe a clear preference for target paper extraction over external literature consultation. Actions targeting the primary paper (“SearchPaper” and “SearchSection”) are invoked 174.1 and 144 times on average, significantly more than “SearchLiterature” which accesses related works only 43.8 times.
5.3 Error Analysis
5.3.1 Syntax Errors
Table A4 shows the syntax error rates for each model across different configurations. Without the Code Agent, syntax errors occurred at rates of 80.3% (“NoAgent”) and 83.3% (“Paper Agent”). After implementing the Code Agent, these error rates dropped significantly to 29.4% (“Code Agent”) and 24.9% (“Sci-Reproducer”). The remaining syntax errors mainly result from incorrectly using repository dependencies. This occurs because our approach, unlike human developers, cannot dynamically access runtime information through a compiler during the code generation process.
5.3.2 Logic Errors
Another issue stems from differences in implementation logic, which can be broadly categorized into: (1) discrepancy in algorithm implementation that result in differing outputs, and (2) missing or mismatch information in the algorithm descriptions in the paper compared to the actual code.
Implementation discrepancy An algorithm may have multiple valid implementation approaches. For example, the cross-entropy loss function can be implemented by directly invoking the PyTorch API “torch.nn.CrossEntropy” or by manually coding it from scratch. Such implementation choices may introduce subtle differences that lead to variations in the final output of the function.
Missing/Mismatched information in algorithm description
Model (Sci-Reproducer) Exe Acc($\uparrow$) CodeBLEU($\uparrow$) RG Acc($\uparrow$) Recall
Intra-File($\uparrow$) Cross-File($\uparrow$) API($\uparrow$)
[Uncaptioned image] GPT-4o-mini 0.220 0.316 0.809 0.588 0.485 0.409
[Uncaptioned image] Deepseek-V3 0.470 0.378 0.834 0.682 0.424 0.609
[Uncaptioned image] o3-mini-low [Uncaptioned image] 0.220 0.292 0.850 0.259 0.091 0.460
Table 4: Experimental Results when missing/mismatched information is regard as external input in the prompt.
Algorithmic descriptions in research papers often lack concrete implementation details, and in certain cases, the provided code may exhibit minor discrepancies compared to the descriptions in the paper. We manually compared the implementation code of all tasks in the dataset with their descriptions in the papers to identify missing or mismatch information. We then provided this information as additional input and apply Sci-Reproducer framework on three LLMs. The Results is shown in Table 4, regarding to Execution Acc, the performance for GPT-4o-mini, Deepseek-V3 and O3-mini-low improved 0.050, 0.250 and 0.040 respectively. The missing information can be divided into four categories:
  • Hyperparameters and configurations: descriptions of target algorithms in papers often omit specific hyperparameter settings, such as the batch size.
  • Numerical stability techniques: standard techniques for ensuring numerical stability, such as handling division by zero.
  • Implementation logic: common implementation practices and model design choices, such as data splitting protocols.
  • Coding strategy: practical programming techniques that enhance implementation efficiency and reliability, such as early stopping criteria.
More examples for each category can be found in Table A5 in Appendix E. As for mismatched information, it occurs far less frequently compared to missing information, and its categories largely overlap with those mentioned above.
To mitigate the widespread issues of missing and mismatched information, the first category can generally be addressed by referencing the original research paper and related literature, or by inspecting the code repository for explicit configurations. However, addressing the other three categories requires familiarity with general machine learning coding conventions, thus necessitating that the LLMs identify and utilize implementation patterns from comparable algorithms to enhance code quality. Future research may improve performance by incorporating implementation insights from similar algorithms through techniques such as in-context learning (Zhou et al., 2023; Xiang et al., 2024), and by leveraging real-time compiler feedback to infer precise variable values.
6 Conclusion
We evaluate LLMs’ ability to replicate algorithms described in recent NLP papers. To support this, we introduce SciReplicate-Bench, a benchmark with rich annotations, and Sci-Reproducer, a dual-agent framework for bridging algorithm understanding and code generation. We assess performance using reasoning graph accuracy and standard implementation metrics. Results show the task is highly challenging, with failures largely caused by missing or inconsistent algorithm descriptions.
7 Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2). We thank the authors of the selected papers for making their code openly available.
Appendix
Appendix A Details of the Annotation Process
Title Conference
1. From Zero to Hero: Cold-Start Anomaly Detection (Reiss et al., 2024) Findings of ACL 2024
2. Addressing Order Sensitivity of In-Context Demonstration Examples in Causal Language Models (Xiang et al., 2024) Findings of ACL 2024
3. Breaking the Ceiling of the LLM Community by Treating Token Generation as a Classification for Ensembling (Yu et al., 2024) Findings of EMNLP 2024
4. Simple but Effective Compound Geometric Operations for Temporal Knowledge Graph Completion (Ying et al., 2024) ACL 2024
5. When is Tree Search Useful for LLM Planning? It Depends on the Discriminator (Chen et al., 2024d) ACL 2024
6. Functional Overlap Reranking for Neural Code Generation (To et al., 2023) Findings of ACL 2024
7. Beyond Single-Event Extraction: Towards Efficient Document-Level Multi-Event Argument Extraction (Liu et al., 2024b) Findings of ACL 2024
8. Unifying Dual-Space Embedding for Entity Alignment via Contrastive Learning (Wang et al., 2024a) COLING 2025
9. Exploring Concept Depth: How Large Language Models Acquire Knowledge and Concepts at Different Layers? (Jin et al., 2024) COLING 2025
10. TRANSMI: A Framework to Create Strong Baselines from Multilingual Pretrained Language Models for Transliterated Data (Liu et al., 2024c) COLING 2025
11. Enhancing Knowledge Distillation of Large Language Models through Efficient Multi-Modal Distribution Alignment (Peng & Zhang, 2024) COLING 2025
12. Document-level Claim Extraction and Decontextualisation for Fact-Checking (Deng et al., 2024) ACL 2024
13. IRCAN: Mitigating Knowledge Conflicts in LLM Generation via Identifying and Reweighting Context-Aware Neurons (Shi et al., 2024) Neurips 2024
14. RouterDC: Query-Based Router by Dual Contrastive Learning for Assembling Large Language Models (Chen et al., 2024b) Neurips 2024
15. Unsupervised Homography Estimation on Multimodal Image Pair via Alternating Optimization (Song et al., 2024) Neurips 2024
16. RAPTOR: Recursive Abstractive Processing for Tree-Organized Retrieval (Sarthi et al., 2024) ICLR 2024
17. Less is KEN: a Universal and Simple Non-Parametric Pruning Algorithm for Large Language Models (Mastromattei & Zanzotto, 2024) Findings of ACL 2024
18. Adaptive Contrastive Search: Uncertainty-Guided Decoding for Open-Ended Text Generation (Arias et al., 2024) Findings of EMNLP 2024
19. MiniCheck: Efficient Fact-Checking of LLMs on Grounding Documents (Tang et al., 2024a) EMNLP 2024
20. Nearest Neighbor Normalization Improves Multimodal Retrieval (Chowdhury et al., 2024) EMNLP 2024
21. Neuron-Level Knowledge Attribution in Large Language Models (Yu & Ananiadou, 2023) EMNLP 2024
22. RaTEScore: A Metric for Radiology Report Generation (Zhao et al., 2024) EMNLP 2024
23. GREEN: Generative Radiology Report Evaluation and Error Notation (Ostmeier et al., 2024) Findings of EMNLP 2024
24. Style-Specific Neurons for Steering LLMs in Text Style Transfer (Lai et al., 2024) EMNLP 2024
25. Language-Specific Neurons: The Key to Multilingual Capabilities in Large Language Models (Tang et al., 2024b) ACL 2024
26. Reasoning Paths Optimization: Learning to Reason and Explore From Diverse Paths (Chia et al., 2024) Findings of EMNLP 2024
27. Lifelong Knowledge Editing for LLMs with Retrieval-Augmented Continuous Prompt Learning (Chen et al., 2024a) EMNLP 2024
28. NeuroMax: Enhancing Neural Topic Modeling via Maximizing Mutual Information and Group Topic Regularization (Pham et al., 2024) Findings of EMNLP 2024
29. Bridging Local Details and Global Context in Text-Attributed Graphs (Wang et al., 2024c) EMNLP 2024
30. Advancing Adversarial Suffix Transfer Learning on Aligned Large Language Models (Liu et al., 2024a) EMNLP 2024
31. SafeDecoding: Defending against Jailbreak Attacks via Safety-Aware Decoding (Xu et al., 2024) ACL 2024
32. MaskLID: Code-Switching Language Identification through Iterative Masking (Kargaran et al., 2024) ACL 2024
33. Towards Robust and Generalized Parameter-Efficient Fine-Tuning for Noisy Label Learning (Kim et al., 2024) ACL 2024
34. Learning to Maximize Mutual Information for Chain-of-Thought Distillation (Chen et al., 2024c) Findings of ACL 2024
35. GLiNER: Generalist Model for Named Entity Recognition using Bidirectional Transformer (Zaratiana et al., 2023) NAACL 2024
36. End-to-End Beam Retrieval for Multi-Hop Question Answering (Zhang et al., 2023a) NAACL 2024
Table A1: List of papers used in our Sci-Replicate benchmark.
Step 1: paper selection
We curated NLP papers from leading conferences in 2024, including ACL, EMNLP, ICLR, Neurips, and COLING. Using a web crawler, we collected accepted paper titles and employed the PapersWithCode API [^5] https://paperswithcode.com/api/v1/docs/ to identify those with open-source implementations. For each identified paper, we retrieved corresponding GitHub repository links and metadata (e.g., stars, issues, release dates) via the GitHub REST API [^6] https://docs.github.com/en/rest?apiVersion=2022-11-28.
To filter candidates, we applied the following criteria:
  • Removed survey/exploratory papers while retaining method-focused research.
  • Applied a cutoff date of January 1, 2024 to avoid data leakage.
  • Excluded repositories with fewer than 5 stars to ensure basic quality assurance.
Subsequently, researchers manually reviewed each candidate paper and its repository. We discarded papers with excessive computational demands, poorly structured code, ambiguous documentation, missing preprocessing steps, or reported reproduction issues.
Step2: python environment setup
For papers passing the initial screening, annotators followed the README to set up the environment and replicate experiments. Common issues included dependency conflicts, data loading failures, and incomplete or buggy code. Annotators attempted to resolve these problems; repositories with irrecoverable errors were excluded.
Step3: annotation
Annotation consists of two steps:
  1. 1.
    Algorithm-Function alignment: most papers contain multiple algorithmic components, often organized as subsections. Annotators segmented these into distinct units and mapped each to its corresponding implementation. Code was refactored to encapsulate each algorithm in a standalone function or method. Papers with implementations too fragmented for restructuring were excluded.
  2. 2.
    Detailed annotation: for each aligned function, annotators documented input/output variables, intra- and cross-file dependencies, and external API usage. Additionally, they inserted explanatory comments mapping code segments to algorithm components. Based on these annotations and variable dependencies, we can construct a reasoning graph representing the implementation logic. During the annotation process, LLMs were employed to assist with algorithm-function alignment and the generation of variable descriptions and code comments. All outputs were subsequently reviewed and corrected by human annotators to ensure accuracy.
The final selected papers are listed in Table A1.
Step 4: verification suite preparation
Finally, annotators created verification suites with 10 test cases per task, drawn from the original datasets used in each repository for the majority of papers. For a small number of repositories, fewer than 10 test cases could be constructed. For instance, algorithms that analyze LLM parameters may have only a single test case. Given the inherent randomness in many NLP implementations and potential machine-related variability, we addressed reproducibility from two angles:
  • Eliminating code randomness: annotators fixed random seeds and replaced non-deterministic operations (e.g., unordered sets) with deterministic equivalents to ensure consistent outputs across runs.
  • Controlling hardware variability: users were instructed to run both reference and generated code locally to eliminate discrepancies caused by hardware differences.
Lastly, annotators implemented task-specific comparison scripts to evaluate output correctness, accounting for variations in return types across tasks.
Appendix B Details of the Task Categories
Figure A1: The categories of the tasks within SciReplicate-Bench.
The benchmark encompasses five main task categories in the NLP domain: representation and embedding methods, loss functions and optimization objectives, information extraction and aggregation, model architecture components, and inference and search algorithms. The distribution of each task category is illustrated in Figure A1.
Appendix C Details of the Actions
In this section, we provide implement details for all actions defined in the Sci-Reproducer.
SearchPaper
We obtain the LaTeX source code of the target academic paper from arXiv [^7] https://arxiv.org/ and apply regular expression-based parsing to extract the content corresponding to each section. Subsequently, we iteratively feed the content of each subsection, along with the query generated by the large language model, into GPT-4o-mini. The model extracts relevant information and returns it as an observation to the paper agent.
SearchSection
Following the same approach as SearchPaper, the tool begins by parsing the LaTeX source code of the target algorithm. Upon receiving a section ID from the Paper Agent, it retrieves and returns the content of the corresponding section.
SearchLiterature
Given a paper ID and a query, the tool attempts to download the corresponding LaTeX source code from arXiv. If the LaTeX source code is unavailable, it returns no information. Otherwise, it extracts content relevant to the query from the paper, following the same procedure as the SearchPaper action.
SearchCode
For each Python file in the code repository, we utilize the Python AST  [^8] https://docs.python.org/3/library/ast.html package to parse the file and extract all defined classes, functions, and global variables. Unlike embedding-based code search methods (Zhang et al., 2024c; 2023c), the Code Agent in our framework directly provides the name of a code item. The tool then returns the corresponding definition if it exists; otherwise, it returns an empty response.
SearchFile
When the Code Agent provides a file name, the tool returns the full content of the corresponding file.
SearchWeb
When the Code Agent issues a query, we use the Google Search API  [^9] https://developers.google.com/custom-search/v1/ to retrieve relevant information from websites. These results are then processed by GPT-4o-mini, which filters the content and extracts the information most relevant to the query for return.
Compiler
Once the Code Agent completes code generation, it invokes the compiler to execute the code. The generated function or method is inserted into the original Python file, and the corresponding Python environment is used to run the code. The output from the compiler is then returned as the feedback.
Appendix D Human Evaluation of Reasoning Graph Accuracy
Table A2: Reasoning Graph Accuracy of Sci-Reproducer Models Evaluated by LLMs and Human Annotators
Sci-Reproducer Model Evaluation Method
GPT-4o Human (Mean ± Std)
GPT-4o-mini 0.782 0.768 (0.066)
O3-mini-medium 0.790 0.789 (0.033)
Table A3: Pearson Correlation between LLM and Human Judgments of Reasoning Graph Accuracy
Number of Data Points r P-value
20 0.7518 <<0.005
To validate the reliability of our LLM-based evaluation metrics, we conducted a human evaluation study on a subset of our benchmark. Three PhD students in computer science independently assessed the reasoning graph accuracy for 20 tasks generated by GPT-4o-mini and O3-mini-medium using Sci-Reproducer.
As shown in Table A2, human evaluations demonstrate strong alignment with our automated LLM-based assessments. The mean scores show consistent agreement between human annotators and GPT-4o evaluations, with standard deviations indicating reasonable inter-annotator consistency.
Furthermore, we computed the Pearson correlation coefficient between human and LLM-based evaluations across all assessed instances. As presented in Table A3, the correlation (r = 0.7518, p ¡ 0.005) indicates a strong positive relationship between human judgments and automated assessments, supporting the validity of our LLM-based evaluation approach.
While this preliminary validation demonstrates the effectiveness of our automated metrics, comprehensive human evaluation across the full benchmark remains an important direction for future work.
Appendix E Figures and Tables
Figure A2: The task consists of two steps: Algorithm Understanding and Code Implementation. (Left) The model must extract an algorithm’s workflow and details from the research paper, including descriptions and variable values from cited papers and other paper sections. (Right) Using this extracted information, the model implements the corresponding function in the code repository, correctly handling dependencies and API calls.
Figure A3: The overview of the SciReplicate-Bench.
Approach Error Ratio ($\downarrow$)
No Agent 80.3
Paper Agent 83.3
Code Agent 29.4
Sci-Reproducer 24.9
Table A4: Syntax error across different settings.
Categories Examples
Hyperparameters and configurations Thresholds, batch sizes, maximum iteration counts, exact numbers of clusters,
initialization methods for variables or vectors, types of regularization (such as L1 or L2)
, and specific distance metrics (e.g., using L2 norm for Euclidean distances)
Numerical stability techniques clamping values to avoid numerical instability, adding small constants during
logarithmic calculations, managing division by zero scenarios, and addressing
rounding and precision issues.
Implementation logic Data splitting, application of dropout, formatting of input sequences,
and handling special or edge cases in the input data.
Coding strategy Caching for performance enhancement, retry mechanisms to handle failures,
early stopping criteria, and strategies for memory optimization.
Table A5: Some examples for different missing information categories.
Figure A4: The prompt for node matching.
Figure A5: The prompt for code generation.
Figure A6: The prompt for Paper Agent.
Figure A7: The prompt for Code Agent.
Figure A8: An example of output report of the Paper Agent.